A certain escalator travels at a rate of 10.6 m/min, and its angle of inclination is 32.5 degrees. What is the vertical component of the velocity? How long will it take a passenger to travel 10.0 m vertically?
Answer by stanbon(75887)

A certain escalator travels at a rate of 10.6 m/min, and its angle of inclination is 32.5 degrees. What is the vertical component of the velocity? How long will it take a passenger to travel 10.0 m vertically?
---
vertical component = 10.6*sin(32.5) = 5.64 m/min
----
time to travel 10m = 10/5.64 = 1.7725 minutes = 1min 46.35 seconds
-----------------
Cheers,
Stan H.
============